OpenStudy (anonymous):

Interval notation of: {x element R : 8 x+3 sqrt(6)<=0 or 8 x>=3 sqrt(6)}

OpenStudy (debbieg):

First you'll need to solve each of the inequalities: \(\Large 8 x+3 \sqrt{6}\le0\ \\ \Large 8 x\ge3 \sqrt{6}\) Since they are joined by an "or", once you know the solution sets to each of the inequalities, you will take the union of those two sets.

OpenStudy (debbieg):

And since they are weak inequalities (less than OR equal to, greater than OR equal to), you will have a [ or ] on the endpoints (except for + or - \(\infty\))